How can I compute and post to journal

The VanDuzer Company has the following footnote to its financial statements:

Note 4: Short-Term Investments

The company holds the following short-term investments at December 31 (in thousands):

Cost Market Value

Trading securities

   U.S. government bonds $670,000 $680,000

Held-to-maturity securities

   Bonds issued by Beta Corp. 540,000 560,000

Available-for-sale securities

   Common shares of Gamma Corp. 600,000 770,000

Compute the amount that VanDuzer would show on its balance sheet for short-term investments.

Suppose the market values of the three securities at the beginning of the year had been as follows (in thousands):

U.S. government bonds $700,000

Bonds issued by Beta Corp. 580,000

Common shares of Gamma Corp. 710,000

Prepare journal entries to recognize the changes in market values that would be recorded in VanDuzer’s books at the end of the year.
Answer #1

1) amount to be computed for short term investment as below:

Difference Actual cost of bond value and market value of securities be posted in books of financial statements by

Trading securities : 670000-680000= 10000

Held to maturities securities: 540000-560000= 20000

Common shares of gamma corporation 600000-770000 = 170000

Overall securities are getting profit: 170000-30000= 140000

Comparing of beginning and ending year for securities as below:

1) trading securities: the difference is 30000 increasing

2)bond issue: the difference is 40000 increasing

3) common shares issues: the difference is 110000 increasing

The overall increase in share by the value of 180000

JOURNAL ENTRIES

Trading securities a/c dr. 10000

To. Cash a/c. 10000

( Being the transaction have the difference between the cost of bond and market value to be recorded)

Bond issue by beta a/c dr 40000

To. Cash a/c. 40000

(Begin shares are issued by the beta company)

Common share a/c dr. 110000

To. Cash. a/c. 110000

(Being shares be issued by gamma corporation)

answered by: WizEdu